# Glassy dynamics in deep neural networks: A structural comparison

### v2

@@ III.3 The long-time limit of $Q(t)$
Figure 4: (a) $Q(t)$ curves collapse onto a master curve for short and intermediate times, when time is rescaled by $\tau_{90}$. Higher temperatures (redder colours) leave the master curve at longer times, and lower values of $Q(t)$. The measurements were made with the same parameters as Fig. 3 and the displayed temperatures are $T=1$, $0.5$, $0.1$, $0.05$, $0.01$, $0.005$, $0.001$, $5\times 10^{-4}$, $10^{-4}$, $10^{-5}$, $10^{-6}$, $10^{-8}$, $10^{-10}$, 00. (b) The final value of the overlap, $Q(t_{f}=3\times 10^{5})$ vs $T$.
@@ III.4 Caging
$\text{MSD}(t,t_{w})=\left\langle\frac{1}{N}\sum_{i=1}^{N}[w_{i}(t+t_{w})-w_{i}(t_{w})]^{2}\right\rangle.$ (7)
As shown in Fig. 5, we see no plateaus, and hence no evidence of caging. In Fig. 5 (a), (b), and (c) the MSD is plotted for regularised, underparameterised, and overparameterised networks respectively, subject to overdamped Langevin dynamics. In all three cases, at high temperatures and intermediate times the MSD demonstrates simple diffusive behaviour with a slope of $1$. In the regularised and underparameterised phases the MSD curves begin to collapse onto a single curve below $T^{\ast}$, i.e., the inflection temperature from the decorrelation plots in Fig. 3. For comparison with the decorrelation plots an MSD equivalent to $\epsilon^{2}$ is shown by the horizontal line. Below $T^{\ast}$ the dynamics are dominated by the gradients of the loss landscape, hence the MSD becomes temperature insensitive, and no longer follows a simple power law in $t$.
Landscape effects are not present in the overparameterised phase as the networks exist in the flat, $\mathcal{L}\approx 0$ plateaus. As such the stochastic part of the dynamics always dominates, and the MSD is always diffusive.
At low temperatures and short times there are oscillations in the MSD in the regularised and underparameterised phases. This is a numerical artefact due to gradients in the loss varying over length scales of a similar size to the learning rate (i.e. step size) used for training. The oscillations disappear when a smaller learning rate is used (not shown), however the long time sections of the curves then become unobtainable.
@@
Figure 5: The MSD of (a) regularised, (b) underparameterised, and (c) overparameterised networks. The dotted black line shows a gradient of 1, i.e. normal diffusion. The solid black line is the MSD corresponding to the value of $\epsilon$ used in Fig. 3. The measurement parameters are the same as those in Fig. 3. The temperatures used in (a) are $T=1$, $0.5$, $0.1$, $0.05$, $0.01$, $0.005$, $0.001$, $5\times 10^{-4}$, $10^{-4}$, $10^{-5}$, $10^{-6}$, $10^{-8}$, $10^{-10}$, 00. In (b) they are $T=1$, $0.5$, $0.1$, $0.05$, $0.01$, $0.005$, $0.001$, $5\times 10^{-4}$, $10^{-4}$, $10^{-5}$, $10^{-6}$, 00. In (c) they are $T=1$, $0.5$, $0.1$, $0.05$, $0.01$, $0.005$, $0.001$, $5\times 10^{-4}$, $10^{-4}$, $10^{-5}$, $10^{-6}$, $10^{-8}$, $10^{-10}$. In all three cases the temperatures run from high (red, top curves) to low (blue, bottom curves). Panel (d) shows the MSD in the regularised phase for networks trained with SGD. The localisation of the networks at long times is visible. The batch sizes used are $b=1$, $2$, $4$, $8$, $16$, $32$, $64$, $100$, $128$, $300$, $600$, $1000$, $3000$, $6000$, $10000$, $60000$. Small batch sizes, corresponding to high noise, are the high curves (red), and large batch sizes are the low (blue) curves.
@@
We hypothesise that the root cause of the lack of caging in DNNs is due to the geometrical complexity of the basins of attraction in the loss landscape. The energy landscape of a structural glass is made up of topologically simple, distinct basins that are roughly spherical in shape as they are formed by the repulsive interactions of the particles acting on length scales of approximately one particle diameter. This situation is depicted schematically in Fig. 6 (left hand panels). These basins have a typical associated length scale, and the barriers separating them a typical height. Consequently the signature plateau of caging is visible in a predictable way when the dynamics are resolved on length scales at least as small as the cage size, and measurements are made over timescales at least as long as the cage-escape time. The situation with the DNN loss landscape is much more complicated. There is no reason for the basins of attraction to have a simple structure, and are likely to be geometrically complex (e.g. being much larger in some dimensions than others) as well as containing topologically non-trivial level sets. The basins have no typical length scale, hence no typical escape time. The recent work of [68] demonstrating the existence of nearly flat paths between distant loss minima, as well as that of [69] mapping the high degree of connectivity between solutions of the spherical negative perceptron, also suggest the basins have a highly extended shape. A schematic loss landscape illustrating these complexities is shown in the right hand panels of Fig. 6, and the real loss landscapes of DNNs are visualised in [70].
@@
What about caging in spin glasses? Here the energy landscape is divided into a high temperature region of saddle points, and a low temperature region of minima, with caging occurring at saddles near the boundary as the number of unstable modes gets very small. Figure 5 suggest this picture does not apply to DNNs and raises the interesting question of whether the topology of the loss and energy landscapes are different, or whether geometric features are sufficient to explain the lack of caging.
Figure 6: Schematics of the energy and loss landscapes of a structural glass and a DNN. The glass energy landscape has simple, approximately spherical basins of a well defined size. In contrast, we expect the DNN loss landscape to have extended basins with no associated length scale.
@@ III.5 The Stokes-Einstein relation and Dynamic Heterogeneity
where the denominator comes from the Stokes drag equation, and the exact constant of proportionality is system dependent [71]. In colloidal glasses (which obey overdamped Langevin dynamics like Eq. (3)) the effective viscosity of the fluid-colloid mixture, $\eta_{\text{eff}}$, is used instead of the bare viscosity of the background fluid [72]. It is a well known signature of the glass transition that the Stokes-Einstein relation begins to be violated as the system enters the supercooled regime [73, 74, 75]. The cause of this violation is thought to be associated with non-Gaussian fluctuations, as measured by the non-Gaussian parameter, $\alpha_{2}$ [76], and the emergence of dynamic heterogeneity (DH). DH is another typical glassy phenomenon, whereby the system begins to separate into spatially localised populations of mobile and non-mobile particles [75, 77, 78]. The mobile particles dominate the diffusion, whereas the non-mobile particles dominate the viscosity, resulting in the decoupling of these two processes, and the violation of Eq. (8).
Figure 7: The non-Gaussian parameter, $\alpha_{2}$, deviates from 0 and hence demonstrates non-Gaussian behaviour in the weight displacements of the regularised (a), underparameterised (b), and overparameterised (c) phases. The measurement parameters are the same as in Fig. 3. The temperatures used in (a) are (going from red, lowest curves, to blue, highest curves) $T=0.1$, $0.01$, $0.001$, $10^{-5}$, $10^{-6}$, $10^{-8}$, $10^{-10}$, 00, in (b) they are $T=0.1$, $0.01$, $0.001$, $10^{-5}$, $10^{-6}$, 00, and in (c) they are $T=0.1$, $0.01$, $0.001$, $10^{-5}$, $10^{-6}$, $10^{-8}$. Histograms of the weight displacements are plotted in the right-hand column, for the regularised (d), underparameterised (e), and overparameterised (f) phases. All phases exhibit bimodal distributions at low temperatures. The temperatures (going from red, right-most curves, to blue, left-most curves) in (d) are $T=10^{-3}$, $10^{-5}$, $10^{-6}$, $10^{-8}$, in (e) they are $T=10^{-2}$, $10^{-3}$, $10^{-5}$, $10^{-6}$, and in (f) they are $T=10^{-1}$, $10^{-3}$, $10^{-6}$, $10^{-8}$.
@@
$\alpha_{2}(t)=\frac{\langle\Delta w^{4}(t)\rangle}{3\langle\Delta w^{2}(t)\rangle^{2}}-1,$ (9)
where $\Delta w$ are the weight displacements [80]. $\alpha_{2}$ is 0 for Gaussian fluctuations in $\Delta w$. As can be seen in Fig. 7 (left-hand column), $\alpha_{2}$ does in fact deviate from 0 for our DNNs, with the deviation starting at long times for low temperatures, and coming to dominate the whole time window as the temperature decreases. This is to be expected as it is the landscape term in Eq. (3) that causes the non-Gaussian behaviour. The $\alpha_{2}$ curves are quite different from those measured in structural glasses, where $\alpha_{2}$ falls again at long times as cage hopping restores Brownian motion [81]. Our data does not reach long enough times to see if this occurs in DNNs, however given the lack of caging we suggest it is unlikely.
We next test for the presence of DH by plotting the distribution of weight motilities, $\Delta w=|w(t_{w}+\Delta t)-w(t_{w})|$, at a sequence of decreasing $T$ values. For structural glasses, a bimodal distribution of mobile and non-mobile particles emerges as $T$ approaches the glass temperature [76]. As shown in Fig. 7 (right-hand column), a bimodal distribution also emerges for DNNs in all three phases. The presence of DH in DNNs is surprising, as DH in structural glasses is typically understood as a result of spatially correlated movements of groups of particles interacting via steric repulsion between neighbours [82, 77, 83]. In contrast, the weights of a DNN interact by complicated, indirect couplings, and their neighbour relations (or equivalently the topology of the space they inhabit) differ considerably from particles packed in 3D Euclidean space.
@@ III.6 Ageing
Ageing, the phenomenon where the behaviour of a system depends explicitly on its age or the waiting time $t_{w}$ (e.g. the time following a quench) is another hallmark of glassy systems. As $T$ decreases towards the glass transition temperature, glass-forming systems fall out of equilibrium, and the resulting broken ergodicity produces ageing effects. The effect of ageing on $Q$ is shown for the regularised and underparameterised phases in Fig. 8 (a) and (c) respectively, as well as the dependence of $\tau_{90}$ on $t_{w}$ in those phases in (b) and (d). Both the regularised and underparameterised regimes exhibit ageing in the landscape-dominated regime $(T<T^{\ast})$, though the effect is stronger in the regularised phase. In both phases a longer $t_{w}$ results in slower dynamics, as it does in glasses. At $T>T^{\ast}$ the networks behave like an equilibrium system and there is no ageing (not shown). The overparameterised phase does not have a transition from noise-dominated to landscape-dominated dynamics as the networks exist in flat plateaus. As such, overparameterised networks never exhibit ageing, as demonstrated in Fig. A11. With the current data it is impossible to make a quantitative statement about the ageing behaviour (e.g. the power law ageing documented in glasses [84, 85, 86]), and varying $t_{w}$ over several more orders of magnitude is computationally unfeasible.
@@ IV Discussion
The presence of DH in DNNs raises the intriguing question of whether there is some “spatial” organisation of the mobile weights, as there is in structural glasses. Such structure could have important practical applications in the sub-field of DNN pruning. Pruning, i.e. reducing a network’s size with minimal reduction in its performance, is an area of growing importance as the immense size of modern networks makes them impractical for use on mobile phones or smart devices, as well as consuming worrying amounts of energy [87, 88, 89]. The presence of two distinct populations of weights, coupled with the insights of disordered physics, could lead to innovative new approaches to pruning.
@@ Appendix A Temperature in networks with overdamped Langevin dynamics
We wish to construct the overdamped Langevin dynamics of a network to be as similar as possible to that of a physical system. The degrees of freedom, the weights, are analogous to the particle positions. The energy of a system is an extrinsic quality, whereas the loss, being an average over the data, is intrinsic. Consequently, the quantity $P_{D}\mathcal{L}$ plays the role of the energy.
@@
Appendix B Choosing $\epsilon$

### v3

@@ III.3 The long-time limit of $Q(t)$
Figure 4: (a) $Q(t)$ curves collapse onto a master curve for short and intermediate times, when time is rescaled by $\tau_{90}$. Higher temperatures (redder colours) leave the master curve at longer times, and lower values of $Q(t)$. The measurements were made with the same parameters as Fig. 3(a) and the displayed temperatures are $T=1$, $0.5$, $0.1$, $0.05$, $0.01$, $0.005$, $0.001$, $5\times 10^{-4}$, $10^{-4}$, $10^{-5}$, $10^{-6}$, $10^{-8}$, $10^{-10}$, 00. (b) The final value of the overlap, $Q(t_{f}=3\times 10^{5})$ vs $T$.
@@ III.4 Caging
$\text{MSD}(t,t_{w})=\left\langle\frac{1}{N}\sum_{i=1}^{N}[w_{i}(t+t_{w})-w_{i}(t_{w})]^{2}\right\rangle,$ (8)
where $t_{w}$ is the waiting time since the system was initialised, as defined for Eq. (5). As shown in Fig. 5, we see no evidence of caging. In Fig. 5 (a), (b), and (c) the MSD is plotted for regularised, underparameterised, and overparameterised networks respectively, subject to overdamped Langevin dynamics. In all three cases, at high temperatures and intermediate times the MSD demonstrates simple diffusive behaviour with a slope of $1$. In the regularised and underparameterised phases the MSD curves begin to collapse onto a single curve below $T^{\ast}$, i.e., the inflection temperature from the decorrelation plots in Fig. 3. For comparison with the decorrelation plots an MSD equivalent to $\epsilon^{2}$ is shown by the horizontal line. Below $T^{\ast}$ the dynamics are dominated by the gradients of the loss landscape, hence the MSD becomes temperature insensitive, and no longer follows a simple power law in $t$.
Landscape effects are not present in the overparameterised phase as the networks exist in the flat, $\mathcal{L}\approx 0$ plateaus. As such the stochastic part of the dynamics always dominates, and the MSD is always diffusive. Furthermore, overparameterised networks trained at $T=0$ undergo almost no motion in weight space, achieving an MSD of around $10^{-16}$ after $10^{5}$ timesteps. We do not include this data in Fig. 5(c) as it falls below floating point precision.
At low temperatures and short times there are oscillations in the MSD in the regularised and underparameterised phases. This is a numerical artefact due to gradients in the loss varying over length scales of a similar size to the learning rate (i.e. step size) used for training. The oscillations disappear when a smaller learning rate is used (not shown), however the long time sections of the curves then become unobtainable. It is also noteworthy that at low temperatures and short times there is a small plateau in Fig. 5(a) and (b), reminiscent of the cage escape plateau. However, the plateau is also present at $T=0$, hence we conclude the plateau is an effect of the landscape dynamics, not thermal noise driving the networks over a barrier.
@@
Figure 5: The MSD of (a) regularised, (b) underparameterised, and (c) overparameterised networks. The dotted black line shows a gradient of 1, i.e. normal diffusion. The solid black line is the MSD corresponding to the value of $\epsilon$ used in Fig. 3. The network and loss parameters are the same as those in Fig. 3(a), (b), and (c) respectively, and $t_{w}=2\times 10^{4}$. The temperatures used in (a) are $T=1$, $0.5$, $0.1$, $0.05$, $0.01$, $0.005$, $0.001$, $5\times 10^{-4}$, $10^{-4}$, $10^{-5}$, $10^{-6}$, $10^{-8}$, $10^{-10}$, 00. In (b) they are $T=1$, $0.5$, $0.1$, $0.05$, $0.01$, $0.005$, $0.001$, $5\times 10^{-4}$, $10^{-4}$, $10^{-5}$, $10^{-6}$, 00. In (c) they are $T=1$, $0.5$, $0.1$, $0.05$, $0.01$, $0.005$, $0.001$, $5\times 10^{-4}$, $10^{-4}$, $10^{-5}$, $10^{-6}$, $10^{-8}$, $10^{-10}$. In all three cases the temperatures run from high (red, top curves) to low (blue, bottom curves). Panel (d) shows the MSD in the regularised phase (same parameters as (a), but $t_{w}=6\times 10^{4}$) for networks trained with SGD. The localisation of the networks at long times is visible. The batch sizes used are $b=1$, $2$, $4$, $8$, $16$, $32$, $64$, $100$, $128$, $300$, $600$, $1000$, $3000$, $6000$, $10000$, $60000$. Small batch sizes, corresponding to high noise, are the high curves (red), and large batch sizes are the low (blue) curves.
@@
We hypothesise that the root cause of the lack of caging in DNNs is due to the geometrical complexity of the basins of attraction in the loss landscape. The energy landscape of a structural glass is made up of topologically simple, distinct basins that are roughly spherical in shape as they are formed by the repulsive interactions of the particles acting on length scales of approximately one particle diameter. These basins have a typical associated length scale, and the barriers separating them a typical height. Consequently the signature plateau of caging is visible in a predictable way when the dynamics are resolved on length scales at least as small as the cage size, and measurements are made over timescales at least as long as the cage-escape time. The situation with the DNN loss landscape is much more complicated. There is no reason for the basins of attraction to have a simple structure, typical length scale, or typical escape time. They are likely to be geometrically complex (e.g. being much larger in some dimensions than others) as well as containing topologically non-trivial level sets. The recent work of [84] demonstrating the existence of nearly flat paths between distant loss minima, as well as that of [85] mapping the high degree of connectivity between solutions of the spherical negative perceptron, also suggest the basins have a highly extended shape.
@@
What about caging in spin glasses? Here the energy landscape is divided into a high temperature region of saddle points, and a low temperature region of minima, with caging occurring at saddles near the boundary as the number of unstable modes gets very small. Figure 5 suggest this picture does not apply to DNNs and raises the interesting question of whether the topology of the DNN loss and glass energy landscapes are different, or whether geometric features are sufficient to explain the lack of caging.
@@ III.5 The Stokes-Einstein relation and Dynamic Heterogeneity
where the denominator comes from the Stokes drag equation, and the exact constant of proportionality is system dependent [86]. In colloidal glasses (which obey overdamped Langevin dynamics like Eq. (3)) the effective viscosity of the fluid-colloid mixture, $\eta_{\text{eff}}$, is used instead of the bare viscosity of the background fluid [87]. It is a well known signature of the glass transition that the Stokes-Einstein relation begins to be violated as the system enters the supercooled regime (the temperature regime where the system beings to fall out of equilibrium, but has not yet cooled sufficiently to be considered a true glass) [88, 89, 90]. The cause of this violation is thought to be associated with non-Gaussian fluctuations of the particles’ displacement distribution, as measured by the non-Gaussian parameter, $\alpha_{2}$ [91, Kob1997a, Donati1998], and the emergence of dynamic heterogeneity (DH). DH is another typical glassy phenomenon, whereby the system begins to separate into spatially localised populations of mobile and non-mobile particles [90, 92, 93]. The mobile particles dominate the diffusion, whereas the non-mobile particles dominate the viscosity, resulting in the decoupling of these two processes, and the violation of Eq. (9).
Figure 6: The non-Gaussian parameter, $\alpha_{2}$, deviates from 0 and hence demonstrates non-Gaussian behaviour in the weight displacements of the regularised (a), underparameterised (b), and overparameterised (c) phases. The measurement parameters are the same as in Fig. 3. The temperatures used in (a) are (going from red, lowest curves, to blue, highest curves) $T=0.1$, $0.01$, $0.001$, $10^{-5}$, $10^{-6}$, $10^{-8}$, $10^{-10}$, 00, in (b) they are $T=0.1$, $0.01$, $0.001$, $10^{-5}$, $10^{-6}$, 00, and in (c) they are $T=0.1$, $0.01$, $0.001$, $10^{-5}$, $10^{-6}$, $10^{-8}$. Histograms of the weight displacements are plotted in the right-hand column, for the regularised (d), underparameterised (e), and overparameterised (f) phases. All phases exhibit bimodal distributions at low temperatures. The temperatures (going from red, right-most curves, to blue, left-most curves) in (d) are $T=10^{-3}$, $10^{-5}$, $10^{-6}$, $10^{-8}$, in (e) they are $T=10^{-2}$, $10^{-3}$, $10^{-5}$, $10^{-6}$, and in (f) they are $T=10^{-1}$, $10^{-3}$, $10^{-6}$, $10^{-8}$. For the distributions in (d), (e), and (f), $\Delta w(t,t_{w})$ was evaluated at $t=2.2\times 10^{5}$. For all plots $t_{w}=2\times 10^{4}$.
@@
$\alpha_{2}(t,t_{w})=\frac{\langle\Delta w^{4}(t,t_{w})\rangle}{3\langle\Delta w^{2}(t,t_{w})\rangle^{2}}-1,$ (10)
where $\Delta w(t,t_{w})=|w(t)-w(t_{w})|$ are the weight displacements [95]. $\alpha_{2}$ is 0 for Gaussian fluctuations in $\Delta w$. As can be seen in Fig. 6 (left-hand column), $\alpha_{2}$ does in fact deviate from 0 for our DNNs, with the deviation starting at long times for high temperatures, and coming to dominate the whole time window as the temperature decreases. This is to be expected as it is the landscape term in Eq. (3) that causes the non-Gaussian behaviour. The $\alpha_{2}$ curves are quite different from those measured in structural glasses, where $\alpha_{2}$ falls again at long times as cage hopping restores Brownian motion [96]. Our data does not reach long enough times to see if this occurs in DNNs, however given the lack of caging we suggest it is unlikely.
We next test for the presence of DH by plotting the distribution of weight displacements, $\Delta w$, at a sequence of decreasing $T$ values. For structural glasses, a bimodal distribution of mobile and non-mobile particles emerges as $T$ approaches the glass temperature [91]. As shown in Fig. 6 (right-hand column), a bimodal distribution also emerges for DNNs in all three phases. The presence of DH in DNNs is surprising, as DH in structural glasses is typically understood as a result of spatially correlated movements of groups of particles interacting via steric repulsion between neighbours [97, 92, 98]. In contrast, the weights of a DNN interact by complicated, indirect couplings, and their neighbour relations (or equivalently the topology of the space they inhabit) differ considerably from particles packed in 3D Euclidean space.
@@ III.6 Ageing
Ageing, the phenomenon where the behaviour of a system depends explicitly on its age or the waiting time $t_{w}$ (e.g. the time following a quench) is another hallmark of glassy systems. As $T$ decreases towards the glass transition temperature, glass-forming systems fall out of equilibrium, and the resulting broken ergodicity produces ageing effects. The effect of ageing on $Q$ is shown for the regularised and underparameterised phases in Fig. 7 (a) and (c) respectively, as well as the dependence of $\tau_{90}$ on $t_{w}$ in those phases in (b) and (d). Both the regularised and underparameterised regimes exhibit ageing in the landscape-dominated regime $(T<T^{\ast})$, though the effect is stronger in the regularised phase. In both phases a longer $t_{w}$ results in slower dynamics, as it does in glasses. At $T>T^{\ast}$ the networks behave like an equilibrium system and there is no ageing (not shown). The overparameterised phase does not have a transition from noise-dominated to landscape-dominated dynamics as the networks exist in flat plateaus. As such, overparameterised networks never exhibit ageing, as demonstrated in Fig. A11. With the current data it is impossible to make a quantitative statement about the ageing behaviour (e.g. the power law ageing documented in glasses [99, 100, 101]), and varying $t_{w}$ over several more orders of magnitude is computationally unfeasible. However, it is interesting to note that the authors of [102] measure sub-aging behaviour in DNNs trained with SGD, where much larger training times are achievable.
@@ IV Discussion
The presence of DH in DNNs raises the intriguing question of whether there is some “spatial” organisation of the mobile weights, as there is in structural glasses. Such structure could have important practical applications in the sub-field of DNN pruning. Pruning, i.e. reducing a network’s size with minimal reduction in its performance, is an area of growing importance as the immense number of parameters in modern networks makes them impractical for use on mobile phones or smart devices, as well as consuming worrying amounts of energy [103, 104, 105]. The presence of two distinct populations of weights, coupled with the insights of disordered physics, could lead to innovative new approaches to pruning.
A challenge inherent in all numerical deep learning investigations, including our own, is the large number of hyperparameters, architectures, minimisation algorithms, and datasets in use across the field. Furthermore, due to the significant computational cost of training large networks, we can only ever hope to sample a small region of this experimental space. For example, we chose not to address finite size effects (e.g. by varying the network and dataset size while keeping $N/P_{D}$ constant) as the present study was already at the computational limits of what can be achieved on a university computing cluster. Where possible we have used physical arguments to choose particular parameters (such as the similarity between the quadratic hinge loss and short-range repulsive interactions), and have provided evidence as to the generality of our findings by repeating them on the CIFAR10 dataset. We hope our work stimulates further research on this topic, extending it to a broader range of deep learning systems.
@@ Appendix A Temperature in networks with overdamped Langevin dynamics
We wish to construct the overdamped Langevin dynamics of a network to be as similar as possible to that of a physical system. The degrees of freedom, the weights, are analogous to the particle positions. The energy of a system is an extensive quality, whereas the loss, being an average over the data, is intensive. Consequently, the quantity $P_{D}\mathcal{L}$ plays the role of the energy.
@@
Appendix B Networks above and below the TTT
In Fig. A1 we measure several characteristics of DNNs above and below the TTT. We train 10 networks at $\lambda=0.04$ (high regularisation, above the TTT), and 10 at $\lambda=0.01$ (low regularisation, below the TTT), initialised with weights randomly drawn from a uniform distribution of width $2/\sqrt{N_{L}}$, where $N_{L}$ is the layer width (this is the default PyTorch initialisation). All networks were trained for 7000 epochs using SGD with a learning rate $\mu=10^{-3}$ and batch size 100, followed by 100 epochs of gradient descent to reach their local minima. In Fig. A1(a) we plot the distribution of weights, demonstrating that above the TTT $w_{i}\approx 0$ $\forall i$, whereas below the TTT, the weights adopt a broad range of values. We plot the distribution of final training loss values above the TTT in Fig. A1(b), and below it in Fig. A1(c). These plots demonstrate that above the TTT $\mathcal{L}\approx 0.5$, as expected for the quadratic hinge loss (Eq. (1)) evaluated at $f=0$, whereas below the TTT the loss adopts a spread of different values. Finally, in Fig. A1(d) and (e), we calculate the Euclidean distance in weight space between all pairs of networks, $|\mathbf{W}_{i}-\mathbf{W}_{j}|$, where $\mathbf{W}_{i}$ is the vector of weights of network $i$. Above the TTT (Fig. A1(d)) all networks exist in the $\mathbf{W}=\mathbf{0}$ state, hence the distance between them is zero. Below the TTT, the networks fall into distinct minima, and the distances between them adopt a spread of values. In the main text we use $\mathcal{L}\approx 0.5$ to indicate that the DNNs are in the quadratic state (i.e. above the TTT), as the loss is a convenient and computationally efficient observable. It is highly unlikely, though not impossible, that an ensemble of networks not in the quadratic state, could also all have $\mathcal{L}\approx 0$.
Figure A1: Networks above and below the TTT. (a) The distribution of weights. (b), (c) The distribution of final training loss values above and below the TTT respectively. (d), (e) The distribution of Euclidean distances between pairs of networks in weight space, above and below the TTT respectively.
Appendix C Choosing $\epsilon$